OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16 x = 21 over 4, x = −11 over 4 x = −1, x = 9 x = −1, x = −9 No solution whats this awnser?

OpenStudy (jdoe0001):

\((\bf -4|x + 5| = -16\implies |x + 5| =\cfrac{-16}{-4}\implies |x + 5| =4\\ \quad \\ |x + 5| =4\implies \begin{cases} +(x + 5) =4\\ \quad \\ \bf -(x + 5) =4 \end{cases}\) so you'd get 2 values for "x", one for each scenario

OpenStudy (anonymous):

whats the awnser a,b,c,d ?@jdoe0001